OpenStudy (lucaz):

\frac{ 10^{15} times sqrt{180} }{ sqrt{320} times \left( 0.01 \right)^{-2} }

OpenStudy (lucaz):

\[\frac{ 10^{15} \times \sqrt{180} }{ \sqrt{320} \times \left( 0.01 \right)^{-2} }\]